A series RLC circuit is connected to a 4.80 kHz oscillator with a peak voltage of 4.00 V. It consists of a 0.300 mH inductor, a 80.0 nF capacitor, and a 35.0 Ω resistor. If ε = ε0, what is the instantaneous current i?
Answer #1

net impedance of circuit

Z^2 = R^2 + ( 2 pi *4800 * 0.3* 10^-3 - 1 / ( 2* 3.14* 4800 * 80* 10^-9))^2

Z = 407.14 9 ohm

peak current is given by

io = Eo / Z = 9.825 mA

=======